OpenStudy (anonymous):

log 7/4=(log7)-(log4) is true or false?

OpenStudy (mertsj):

What is log7-log4?

OpenStudy (anonymous):

wat do you mean?

OpenStudy (mertsj):

Have you had this law of logs: \[\log (m)-\log(n)=\log \frac{m}{n}\] ?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Then what is log 7 - log 4?

OpenStudy (anonymous):

log7/4

OpenStudy (mertsj):

So is that statement you posted true or false?

OpenStudy (anonymous):

true

OpenStudy (mertsj):

Very good.

OpenStudy (anonymous):

Thanks:)

OpenStudy (mertsj):

yw